OpenStudy (marmar10):

How many moles of nitrogen, N, are in 73.0 g of nitrous oxide, N2O?

OpenStudy (anonymous):

\[73.0g N2O \times \frac{ 1mol N _{2} O }{ 44g N _{2} O } \times \frac{ 2mol N }{ 1mol N _{2} O }\]